OpenStudy (anonymous):

0.21212121212121......... = 7/33?

OpenStudy (kamille):

Yes, you are right. Because : as you know 0.21212121212121 is the same as 0.(21), right? so if we say that 0.(21)=x|*100 Now you get 21,(21)=100x and finally 100x-x=21,(21)-0.(21) is 99x=21 |:99 and you get 21/99 and it is the same as 7/33

OpenStudy (radar):

dividing 7 by 33 gives you .21212121212121 It would then be safe to say, "Yes"